Lease Terms
• Rental Rate: $40 NNN for five years and then $45NNN until the end of term
• Term- 10 years
• Size: 10,000 sf
• Free Rent- 12 months free (outside term, net free rent)
• Tenant Improvement- $70 sf
• Commission – 4% for five years, 2% thereafter
• Operating Expenses $15.00 (3% growth)
• Amortization Rate 8%
• Calculate the following:
(1) Average Rent
(2) Effective Rent
(3) Commission
(4) NPV/sf
(5) NER
(6) What is the NER?
(7) What is the NER if it was 12 months GROSS free?
Answer #1

Answer 1) average rent = (40*5+45*5)/10 = 42.5$*10000 = 425000$

Answer 2) effective rent = (40*5+45*5)/10+1 = 386300$ (1 month free rent)

Answer 3) commission = (4%*5*40+2%*5*45)*10000 = 125000$

Answer 4) npv/sf => pv of 8% for 10 years is 6.71 approx

42.5*6.71 = 285.175 - (15*11.46*0.46) [11.46 is the fv annuity by 3% on operating expense and 0.46 is pv factor by 8%]

285.175 - 79.075 = 206.1

For npv/sf 206.1 - 70 = 146.1 (answer4)

Answer 5) NER = effective rate - expenses

38.63-15 = 23.63 /sf

Answer 6) net effective rate means the net calculation of the rent recieved after deducting the incurred expenses

Answer 7) in that case effective rate will be 42.5*11/12 = 38.95/sf and ner = 38.95 - 15 = 23.95/sf